OpenStudy (anonymous):

Solve 0.25(d-12) = 4

OpenStudy (anonymous):

0.25d - 3 = 4 0.25d = 7 d = 28

OpenStudy (anonymous):

0.25(d-12)=4 multiply all terms times 4: d-12=16 add 12 to both sides of the equation: d=28